# Stuffed Mamba: State Collapse and State Capacity of RNN-Based Long-Context Modeling

### v1

@@ Abstract
One essential advantage of recurrent neural networks (RNNs) over transformer-based language models is their linear computational complexity concerning the sequence length, which makes them much faster in handling long sequences during inference. However, most publicly available RNNs (e.g., Mamba and RWKV) are trained on sequences with less than 10K tokens, and their effectiveness in longer contexts remains largely unsatisfying so far. In this paper, we study the cause of the inability to process long context for RNNs and suggest critical mitigations. We examine two practical concerns when applying state-of-the-art RNNs to long contexts: (1) the inability to extrapolate to inputs longer than the training length and (2) the upper bound of memory capacity. Addressing the first concern, we first investigate state collapse (SC), a phenomenon that causes severe performance degradation on sequence lengths not encountered during training. With controlled experiments, we attribute this to overfitting due to the recurrent state being overparameterized for the training length. For the second concern, we train a series of Mamba-2 models on long documents to empirically estimate the recurrent state capacity in language modeling and passkey retrieval. Then, three SC mitigation methods are proposed to improve Mamba-2’s length generalizability, allowing the model to process more than 1M tokens without SC. We also find that the recurrent state capacity in passkey retrieval scales exponentially to the state size, and we empirically train a Mamba-2 370M with near-perfect passkey retrieval accuracy on 256K context length. This suggests a promising future for RNN-based long-context modeling.
footnotetext: Corresponding authors
@@ 1 Introduction
Recent transformer-based language models (Vaswani et al., 2017; Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023; Dubey et al., 2024) have demonstrated impressive capabilities in reasoning over long sequences with thousands and even millions of tokens (Team, 2024a; b; GLM et al., 2024). However, they rely on the attention mechanism that scales quadratically regarding the sequence length, making them extremely costly for inference over long sequences. In contrast, recurrent neural networks (RNNs) (Bengio et al., 1994) have a contextual memory with constant state size. Thus, during inference, their per-token computational and memory complexity scales linearly with the sequence length, making them much more efficient in processing long sequences.
Despite the promising future of RNNs to process long contexts in terms of efficiency, their long-context performances are far from satisfying. Most recent state-of-the-art (SOTA) RNN-based language models (hereafter referred to as RNNs for simplicity), such as Mamba-1 (Gu & Dao, 2023), Mamba-2 (Dao & Gu, 2024), the RWKV series (Peng et al., 2023a; 2024), and GLA (Yang et al., 2023) are trained on sequences with less than 10K tokens. Existing works have shown that Mamba-1 and RWKV-4 suffer from severe performance drops when the context length exceeds their training length (Ben-Kish et al., 2024; Zhang et al., 2024a; Waleffe et al., 2024).
In this paper, we study the problem of what causes the current RNNs’ inability to handle long contexts and the possible solutions for supporting long contexts. When applying RNNs to longer contexts, we observe two critical problems. (1) RNNs are unable to generalize along the sequence length. They exhibit abnormal behavior when the context length exceeds the training length, resulting in poor long-context performance. (2) Since their memory size is constant, although they can process infinitely long inputs, there is an upper bound to the amount of information the state can represent. Therefore, there is an upper bound of the contextual memory capacity (the maximum number of tokens that can be remembered), and tokens beyond that limit will be forgotten.
Then we dive deeper into the formation of the above problems. We first attribute the cause of the length generalization failure of SOTA RNNs to a phenomenon we call state collapse (SC). We inspect the memory state distribution over time and discover that its collapse is caused by a few dominant outlier channels with exploding values. These outliers cause vanishing values in other channels when the output hidden representation is normalized. By analyzing various components of the state update rule, we show that SC is caused by the inability to forget the earliest tokens (by decaying the state with a smaller multiplier) when there is more information than it can remember.
Underpinned by these analyses, we propose three training-free techniques for mitigating the collapse and one mitigation method based on continual training on longer sequences. Our methods rely on forcing the model to forget contextual information by reducing the memory retention and insertion strength, normalizing the recurrent state, or reformulating the recurrence into an equivalent sliding window state.
Empirical results on Mamba-2 reveal that our training-free SC mitigation methods allow the model to consume more than 1M tokens without SC. By further training on longer sequences that exceed the model’s state capacity, we empirically verify that for a given state size, there is a threshold for training length beyond which the model will not exhibit SC. This insight allows us to establish a relationship between state capacity and state size. Then, by training Mamba-2 models of different sizes, we establish that the capacity is a linear function of the state size. Furthermore, we conduct the same experiments on the widely used passkey retrieval task (Mohtashami & Jaggi, 2023), and show that the length where Mamba-2 has near-perfect passkey retrieval accuracy is an exponential function of the state size. The experiment results in a Mamba-2 370M model that can achieve near-perfect passkey retrieval accuracy on 256K context length, significantly outperforming transformer-based models of the same size in both retrieval accuracy and length generalizability. Our results show that the commonly used training lengths for RNN-based models may be suboptimal and that RNN-based long-context modeling has promising potential.
Our contributions can be summarized as follows.
  • We provide the first systematic study regarding state collapse, a phenomenon that results in the length generalization failure of RNNs.
  • We propose three training-free mitigation methods and one method based on continual training to improve the length generalizability of RNNs.
  • By analyzing the hidden representations, we attribute state collapse to state overparameterization, which establishes the connection to state capacity. Based on this analysis, we empirically estimate the state capacity of Mamba-2 as a function of the state size.
  • We train and release the first RNN-based language model with near-perfect accuracy on passkey retrieval with 256K tokens. It has only 370M parameters, making it the smallest model with near-perfect passkey retrieval accuracy at this length at the time of writing.
[…]
(b) Mamba-2 series on long documents from RedPajama.
(c) RWKV-6 1.6B and 3B on the “newlines” prompt.
Figure 1: The perplexity of Mamba-2 and RWKV-6 as a function of token position on real and synthetic data. The red dotted line represents the models’ training lengths. They all fail to extrapolate.
Model checkpoints and source code are released at https://www.github.com/thunlp/stuffed-mamba.
2 Related Works
RNN-Based Language Models
There is a recent surge of interest in RNN-based language models, because, contrary to transformer-based ones, their per-token inference cost does not increase with the sequence length. Linear Attention (Katharopoulos et al., 2020) replaces the softmax attention in transformer-based models with kernel-based approximations that have equivalent recurrent formulations. Some notable recent RNNs include the RWKV series (Peng et al., 2023a; 2024), the Mamba series (Gu & Dao, 2023; Dao & Gu, 2024), Gated Linear Attention (Yang et al., 2023), among others (Zhang et al., 2024b; Yang et al., 2024; De et al., 2024; Arora et al., 2024; Orvieto et al., 2023; Sun et al., 2023). These models have shown strong capabilities in many language processing tasks, sometimes outperforming transformer-based models. However, as we will empirically show, some of these models fail to extrapolate much beyond their training length.
Some transformer-based models have adopted sliding window attention (Beltagy et al., 2020; Jiang et al., 2023), which essentially turns them into RNNs. However, these models have been shown to perform poorly in long-context tasks and fail to extrapolate to very long contexts (Zhang et al., 2024a).
Length Generalization
Most SOTA language models in the last few years have been based on the transformer (Vaswani et al., 2017) architecture. These models, when using certain variants of position encoding, can process arbitrarily long sequences. However, they exhibit severe performance drops on tokens beyond the training length (Zhao et al., 2024). To alleviate this shortcoming, many works have focused on modifying positional encoding (Peng et al., 2023b; Zhu et al., 2023; Ding et al., 2024; Jin et al., 2024), some achieving training-free length generalization to certain extents. Similarly in spirit, this study also explores some post-training modifications for enhancing RNN-based models’ length generalization capabilities.
Length Generalization of Mamba
Some concurrent works have explored extending Mamba’s context length by controlling the discretization term ($\Delta_{t}$ in Eq. 3) (Ben-Kish et al., 2024), such as dividing it by a constant to make it smaller [^1] https://github.com/jzhang38/LongMamba. This essentially makes the memory decay factor ($\alpha_{t}$ in Eq. 4) closer to 1, which makes the state retain more contextual information. However, it also unnecessarily diminishes the inserted information on all tokens. Consequently, although it can mitigate SC, it results in poor performance in the passkey retrieval task (details in Appendix C).
3 Preliminary
Mamba-2
The Mamba-2 architecture consists of a stack of $L$ Mamba-2 layers, each Mamba-2 layer consists of $H$ heads that are computed in parallel, and the output of a layer is the sum of the output of each head. Each head in each layer can be formulated as follows. [^2] We use row-vector representation, so $a^{T}b$ denotes an outer product of two vectors.
$y_{t}$ $=\text{Norm}(o_{t}\odot u_{t}W_{\text{gate}})W_{o}\in\mathbb{R}^{d}$ (1)
$o_{t}$ $=C_{t}h_{t}+D\odot x_{t}\in\mathbb{R}^{P}$ (2)
$h_{t}$ $=h_{t-1}\overline{A}_{t}+\overline{B}_{t}^{T}x_{t}\in\mathbb{R}^{N\times P}$ (3)
where $t$ denotes the current time step, $u_{t},y_{t}\in\mathbb{R}^{d}$ are the input and output hidden representations of the $t$-th token, $\text{Norm}(\cdot)$ denotes the RMS normalization (Zhang & Sennrich, 2019), $D\in\mathbb{R}^{P}$, $W_{\text{gate}}\in\mathbb{R}^{d\times P}$, and $W_{o}\in\mathbb{R}^{P\times d}$ are trainable parameters, $\odot$ denotes element-wise product, and $d,N,P$ are hyperparameters, denoting the hidden dimensionality, state dimension, and head dimension, respectively. The other variables are parameterized as follows:
$\overline{A}_{t}$ $=\alpha_{t}I\in\mathbb{R}^{P\times P}$ (4)
$\overline{B}_{t}$ $=\Delta_{t}B_{t}\in\mathbb{R}^{N}$ (5)
$\alpha_{t}$ $=\exp(-\Delta_{t}\exp(A))\in\mathbb{R}$ (6)
$C_{t}$ $=\sigma(\text{Conv}(u_{t}W_{C}))\in\mathbb{R}^{N}$ (7)
$B_{t}$ $=\sigma(\text{Conv}(u_{t}W_{B}))\in\mathbb{R}^{N}$ (8)
$x_{t}$ $=\sigma(\text{Conv}(u_{t}W_{x}))\in\mathbb{R}^{P}$ (9)
$\Delta_{t}$ $=\text{Softplus}(u_{t}W_{\Delta}+b_{\Delta})\in\mathbb{R}$ (10)
where $\left(W_{C},W_{B}\in\mathbb{R}^{d\times N},W_{x}\in\mathbb{R}^{d\times P},W_{\Delta}\in\mathbb{R}^{d\times 1},b_{\Delta},A\in\mathbb{R}\right)$ are trainable model parameters. $\text{Conv}(\cdot)$ denotes a channel-wise one-dimensional convolutional layer (more details in Appendix A.2). $\sigma$ denotes the SiLU function (Elfwing et al., 2017). Importantly, $h_{t}$ is called the hidden state or recurrent state, which is contextual memory that stores information from all tokens up to $t$. $\alpha_{t}$ is the decay multiplier that controls the strength of memory decay (i.e., forgetting). A more detailed formulation is given in Appendix A.
It is worth mentioning that since $\overline{A}_{t}$ is scalar-identity, the update rule can be seen as a variant of Gated Linear Attention (Yang et al., 2023) and is very similar to many existing RNNs such as RWKV (Peng et al., 2024) and RetNet (Sun et al., 2023). Thus, much of the experiments and insights may apply to other architectures. We leave such exhaustive ablation studies for future work.
4 State Collapse
We first examine state collapse (SC)—a phenomenon that causes RNN models to exhibit abnormal behaviors on inputs longer than those seen during training. We analyze the effects of SC on language modelings and the passkey retrieval task. Then we trace SC to the components of the state’s update rule and provide an explanation regarding training length overfitting. Finally, we propose three training-free mitigation methods by modifying the update rule and one method based on continual pre-training on longer sequences to avoid overfitting.
4.1 Length Generalization Failure
Language Modeling
[…]
Passkey Retrieval Evaluation
Language modeling may not reflect downstream capabilities, thus, we evaluate several strong RNNs on the passkey retrieval task (Mohtashami & Jaggi, 2023; Zhang et al., 2024a), a simple synthetic task where a model is prompted to recall a 5-digit passkey from a lengthy context. The hyperparameters and results of other RNNs can be found in Appendix B and D. The results for Mamba-2 are reported in Figure 2. We find that Mamba-2 models fail to generalize to sequences longer than the training length. For instance, Mamba-2, trained on 8K context windows, has near-perfect retrieval accuracy within 8K contexts (except for the smaller 130M checkpoint), but poor or even zero accuracy on sequences longer than 16K, regardless of model sizes.
@@ 4.1 Length Generalization Failure
Figure 2: The performance of Mamba-2 official checkpoints on the passkey retrieval task.
(a) 130M
(b) 370M
(c) 780M
(d) 1.3B
Figure 3: The accuracy of Mamba-2 official checkpoints on the passkey retrieval task where the answer is in the last $n$ tokens, with $n$=1K, 2K, 4K, and 8K.
This behavior is unexpected because the update rule (Eq. 3) has a stable exponential memory decay (it converges to a constant value if the variables are fixed). Therefore, we expect that RNNs of such form should have a good retrieval accuracy on the last $k$ tokens, and tokens earlier than that are forgotten. This unexpected finding also implies that when processing contexts longer than the training length $T_{\text{train}}$, it is better to keep just the last $T_{\text{train}}$ tokens and discard everything else. However, this is not trivial in an online inference scenario because all token information is compressed into a single state.
4.2 What is the Cause of State Collapse?
Figure 4: The mean and variance of the first 8 heads in the 38th layer of Mamba-2 370M. It exhibits a clear explosion when $t$ is greater than the training length. The red dotted line indicates the training length.
Figure 5: The distribution of the channels in a collapsing state (head 2 for the 38th layer) at two different time steps.
Since the recurrent state’s dimensionality does not change over time, the sharp change of behavior during state collapsing must be a result of a change in the state’s value. We inspect the statistics of the recurrent states of each layer in Mamba-2 370M and find that the mean and variance of some heads change sharply when the context length exceeds the training length, as shown in Figure 5. Appendix G includes a more detailed report on the statistics of each head. The state at $t=20K$ of one head with exploding variance is shown in Figure 5. From it, we discover that this variance explosion can be largely attributed to a few outlier channels, while most channels are relatively stable.
We emphasize that SC occurs largely independent of the prompt, occurring in both pre-training data samples and generated meaningless texts, even for prompts consisting of only whitespace characters (Figure 1 (a) and (b) shows SC on two different prompts). This means that the information inserted into the state does not cause its collapse.
To further attribute to the specific variables that cause SC, we inspect the values of $\Delta_{t}$, $B_{t}$, and $x_{t}$ on various heads with collapsing states. Figure 6 reports one example of the inspection, we can see that $x_{t}$ is relatively stable compared to the $\Delta_{t}$ and $B_{t}$, even though they are all functions of $u_{t}$. We also notice that $B_{t}$ explodes earlier than $\Delta_{t}$. Therefore, we conclude that the collapse is largely attributable to $B_{t}$. Further inspection reveals that the convolutional weights that generate $\Delta_{t}$ and $B_{t}$ (Eq. 8 and 10) are noticeably greater in variance than those for $x_{t}$ (Eq. 9). We leave a more in-depth attribution study for future work.
(a) $\Delta_{t}$. Each curve represents a head.
(b) $B_{t}$. Each curve represents a channel.

### v2

# Stuffed Mamba: Oversized States Lead to the Inability to Forget

@@ Abstract
Recent advancements in recurrent architectures, such as Mamba and RWKV, have showcased strong language capabilities. Unlike transformer-based models, these architectures encode all contextual information into a fixed-size state, leading to great inference efficiency. However, this approach can cause information interference, where different token data conflicts, resulting in performance degradation and incoherent outputs beyond a certain context length. To prevent this, most RNNs incorporate mechanisms designed to “forget” earlier tokens. In this paper, we reveal that Mamba-based models struggle to effectively forget earlier tokens even with built-in forgetting mechanisms. We demonstrate that this issue stems from training on contexts that are too short for the state size, enabling the model to perform well without needing to learn how to forget. Then, we show that the minimum training length required for the model to learn forgetting scales linearly with the state size, and the maximum context length for accurate retrieval of a 5-digit passkey scales exponentially with the state size, indicating that the model retains some information beyond the point where forgetting begins. These findings highlight a critical limitation in current RNN architectures and provide valuable insights for improving long-context modeling. Our work suggests that future RNN designs must account for the interplay between state size, training length, and forgetting mechanisms to achieve robust performance in long-context tasks. [^1] Code is released at https://www.github.com/THUNLP/stuffed-mamba
@@ 1 Introduction
Transformer-based large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024) have shown impressive capabilities in processing very long sequences (Gemini Team et al., 2024; MiniMax et al., 2025). However, these models incorporate self-attention (Vaswani et al., 2017) whose complexity scales quadratically with sequence length, making long-context processing costly. In contrast, recurrent neural networks (RNNs) (Bengio et al., 1994) have a fixed-size contextual memory. Thus, their per-token time and space complexities are constant and they are much more efficient for long sequences. Despite this advantage, their effectiveness in modeling long contexts remains underexplored. Most recent state-of-the-art (SOTA) RNNs, such as Mamba-1 and Mamba-2 (Gu & Dao, 2023; Dao & Gu, 2024), GLA (Yang et al., 2024a), and RWKV (Peng et al., 2024a) are trained on context lengths below 10K, and existing works have shown that their performance degrades sharply when the context length exceeds the model’s training length [^2] Throughout this paper, “training length” refers to the context length used during training. (Ben-Kish et al., 2024; Zhang et al., 2024a; Waleffe et al., 2024).
Figure 1: The LM loss of Mamba-2 as a function of token position. The training length is 8K.
@@
Figure 2: The accuracy of Mamba-2 on the passkey retrieval task. “Ans. Depth” refers to the passkey position divided by the context length.
In this paper, we analyze the factors that cause the inability of Mamba-based models to handle contexts longer than the training length. By inspecting memory retention strength and modifying the forgetting mechanism, we discover that the performance drop is caused by the inability to forget earlier tokens, although Mamba has a built-in forgetting mechanism. Insufficient forgetting leads to interference between multiple token representations, causing faulty memory recall and, ultimately, performance degradation in longer contexts. We provide two lines of evidence for this discovery: (1) The first token’s retention strength is very high throughout its training context window. (2) Artificially inducing forgetting via interventions of the state’s update rule can mitigate this performance degradation.
We hypothesize that the inability to learn an effective forgetting mechanism is due to state overparameterization—where the model’s state is excessively large, allowing it to minimize language modeling loss without much forgetting. Two key pieces of evidence support this hypothesis. (1) Initially, the model demonstrates robust forgetting, retaining only the last $k$ tokens and forgetting earlier ones. However, as training progresses, the model’s ability to forget diminishes while its recall of contextual information improves, resembling overfitting. This suggests that the model increasingly attempts to retain all available information within the context. (2) We observe that forgetting occurs only when the training context length exceeds the state’s capacity to retain all information, forcing the model to forget less relevant details. Notably, larger states require longer training context lengths to effectively learn and implement forgetting.
We next investigate the minimum training length required for Mamba-2 to effectively learn forgetting and the maximum context length in which the model can recall information. First, by varying model sizes and training lengths, we observe that the training length threshold scales linearly with the state size, confirming that forgetting only occurs when the training length exceeds the model’s state capacity. Second, while this threshold represents the point at which the contextual information exceeds the state’s capacity, we demonstrate that the model can still recall tokens beyond this context window. Evaluation on passkey retrieval (Mohtashami & Jaggi, 2023)—a simple retrieval task—shows that the maximum context length with perfect retrieval accuracy scales exponentially with state size. Notably, with continued pre-training, Mamba-2 with 370M parameters achieves near-perfect retrieval on a 256K context length, outperforming similarly sized transformer models. These findings suggest that current training lengths for RNN models may be suboptimal and underscore the potential of RNN-based architectures for modeling long-context sequences.
This paper is structured as follows. Section 2 describes the Mamba-2 architecture and provides evaluation results showing its inability to generalize beyond its training length. Section 3 provides arguments for the importance of forgetting and provides evidence showing that the model has failed to learn a robust forgetting mechanism. Section 4 presents a high-level explanation for why Mamba-2 has failed to learn how to forget. Finally, in Section 5, experiments are conducted to verify the claims and provide important conclusions for training long-context recurrent models.
The main findings of this paper can be summarized as follows:
The inability to forget. We discover that Mamba-2, and RWKV-6 (some SOTA RNNs) do not know how to robustly forget earlier information to avoid memory overload. This causes performance degradation for contexts longer than the training length.
State overparameterization. We provide overparameterization as a plausible explanation for the inability to forget and provide empirical evidence for this hypothesis.
Minimum training length of forgetting. In alignment with the state overparameterization hypothesis, we empirically discover that for any state size, there exists a training length threshold where the Mamba-2 learns forgetting if and only if the training length is above that threshold. We also find that this relationship is linear.
2 Preliminaries
In this section, we first describe the Mamba-2 architecture and corresponding notations. Then, we evaluate Mamba-2 on language modeling and passkey retrieval with context lengths exceeding their training length to illustrate the consequence of the inability to forget.
2.1 Mamba-2
The Mamba-2 architecture consists of $L$ layers, each consisting of $H$ heads computed in parallel. The layer’s output is the sum of the heads’ outputs. Let $u_{t}\in\mathbb{R}^{d},y_{t}\in\mathbb{R}^{P}$ denote the input and output vectors of the layer at $t$ time step. The computation at $t$ time step for each head can be formulated as follows:
$y_{t}$ $=C_{t}h_{t}\in\mathbb{R}^{1\times P}$ (Query rule) (1)
$h_{t}$ $=h_{t-1}\underbrace{\alpha_{t}}_{\text{Decay}}\underbrace{+\overline{B}_{t}x_{t}}_{\text{Insertion}}\in\mathbb{R}^{N\times P}$ (Update rule) (2)
where $C_{t}\in\mathbb{R}^{P\times N},\overline{B}_{t}\in\mathbb{R}^{N\times 1},x_{t}\in\mathbb{R}^{1\times P},\alpha_{t}\in\mathbb{R}$ are functions of $u_{t}$, $d,N,P$ are hyperparameters, denoting the hidden dimensionality, state dimension, and head dimension, respectively, and $h_{t}$ is the $t$-th recurrent state. Eq 1 and 2 are called the “query rule” and “update rule” because they determine how memory is queried from the recurrent state, and how the state is updated.
The other variables are parameterized as follows:
$\overline{B}_{t}$ $=B_{t}\Delta_{t}\in\mathbb{R}^{N\times 1}$ (3)
$\alpha_{t}$ $=\exp(-\Delta_{t}\exp(A))\in\mathbb{R}$ (4)
$\Delta_{t}$ $=\text{Softplus}(u_{t}W_{\Delta}+b_{\Delta})\in\mathbb{R}$ (5)
where $B_{t}\in\mathbb{R}^{N\times 1}$ is a function of $u_{t}$ and $A,b_{\Delta}\in\mathbb{R},W_{\Delta}\in\mathbb{R}^{d\times 1}$ are trainable model parameters. Appendix A presents more details on the model. Notably, Mamba-2’s update rule is similar to many existing RNNs (Peng et al., 2024a; Sun et al., 2023; Yang et al., 2024a). Thus, some conclusions/insights may apply to other architectures. We leave such exhaustive ablation studies for future work.
Importantly, $h_{t}$ is the contextual memory representation that stores information from all tokens up to $t$. $\alpha_{t}\in(0,1)$ is the memory decay multiplier that controls the strength of forgetting. Past information is completely forgotten when $\alpha_{t}\rightarrow 0$ and completely retained with $\alpha_{t}\rightarrow 1$. In this paper, we refer to $\alpha_{t}$ as the memory retention strength.
2.2 Length Generalization Failure of Mamba-2
Language Modeling
Figure 2 shows the language modeling loss of Mamba-2 as a function of token position. The result shows that Mamba-2 models suffer great performance degradation when the context length is much longer than their training lengths. Furthermore, we find that larger models have worse length generalization abilities.
Passkey Retrieval Evaluation
Language modeling may not reflect downstream capabilities, thus, we also evaluate Mamba-2 recall ability on the passkey retrieval task (Mohtashami & Jaggi, 2023; Zhang et al., 2024a). It is a widely-used simple synthetic task where a model is prompted to recall a 5-digit passkey from a lengthy context.
The passkey retrieval result is reported in Figure 2. We find that Mamba-2 (except for the smaller 130M checkpoint) has near-perfect retrieval accuracy within 8K tokens, but poor or even zero accuracy on sequences longer than 16K, regardless of model sizes. This behavior is unexpected because the update rule (Eq. 2) has a stable exponential memory decay (it converges to a constant value if the variables are fixed). Therefore, we expect RNNs of such form to have a good retrieval accuracy on the last $k$ tokens, and tokens earlier than that are forgotten. However, when the context is too long, Mamba-2 fails to even recall very recent tokens. This implies that the limitation is not the inability to retain memory about the answer, but the inability to forget irrelevant past tokens.
More experimental details and the results of some other recurrent architectures can be found in Appendix B and D.
3 The Inability to Forget
In this section, we argue that Mamba-2’s length generalization failure can be attributed to its inability to forget contextual information. We first provide arguments for the importance of forgetting. Then, we present empirical evidence to verify that the model has not learned a robust forgetting mechanism. Furthermore, we show how the inability to forget is manifested in the statistics of the recurrent state.
3.1 The Importance of Forgetting
The fact that the model fails to retrieve tokens at any position when the context length exceeds a certain threshold has one critical implication: the existence of earlier tokens damages the model’s ability to recall both earlier and more recent tokens. This is because the model’s state $h_{t}$ at time step $t$ can be formulated as a weighted sum of past information:
When we try to retrieve the memory inserted at time step $s$, we would query the state with $C_{t}=\overline{B}_{s}$, which returns:
$y_{t}=C_{t}\sum_{i=1}^{t}\alpha_{i:t}\overline{B}_{i}x_{i}=\alpha_{s:t}(C_{t}\overline{B}_{s})x_{s}+\underbrace{\sum_{i\neq t}\alpha_{i:t}C_{t}\overline{B}_{i}x_{i}}_{\text{Retrieval error}},$ (7)
When all $\overline{B}_{i}$ are mutually orthogonal, querying the memory with $C_{t}=\overline{B}_{s}$ returns a scaled version of $x_{s}$. However, as the context length increases, $B_{s}$ cannot be mutually orthogonal, in which case multiple memory entries interfere and cause retrieval errors. A small error may not affect the final output because subsequent calculations may tolerate these errors. However, if there are many preceding tokens, then this retrieval error can be too large. To optimize retrieval accuracy, the model needs to produce a small enough decay $\alpha_{t}$ to diminish the inference from earlier tokens, which can be viewed as forgetting them.